Short Report – QA/QC Meeting May 2018

Date: 2018-05-17

Participants / Excused

LH, IB, PM, MG, SB, WE, TB, RM, IF, KF, Kritika (internship)

Talking Points


  1. If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
  2. SRP & Tech-SRP: please send short statement about site to LH (max. 2 sentences)Automatic
  3. Automatic fluxes are active again and calculated each day for the last x days (currently last 7 days).
  4. Online QA/QC checks started for CH-DAV and CH-FRU: Link Next step is to implement this check for CH-CHA and CH-OE2, please contact LH.
  5. All calculations, e.g. QuickViews, fluxes, will be implemented on the grassland server  due to ongoing RDS problems. Last flux calculations worked for CH-DAV. For now, you can try to run it on the RDS, it seems to work sometimes.
  6. Preview Swiss FluxNet homepage:
  7. Regarding GIN: RSS feed needed so that all news from GIN are automatically displayed on Swiss FluxNet homepage.
  8. New Graphical Meteo Screening Tool by JE:


  1. [DONE] Automatic fluxes will be activated again (goal is April 2018), priority conflicts w/ ICOS. All sites now active.
  2. [IN PROGRESS] SRP QA/QC Weekly checks in online checklists (instead of the server Excel sheet), CH-DAV and CH-FRU started. Next, CH-CHA and CH-OE2, please contact LH.
  3. [?UPCOMING?] Tech-SRP QA/QC Weekly checks in online checklists. As soon as ICOS allows.
  4. [UPCOMING] On-demand flux calculations on grassland server instead of RDS. As soon as ICOS allows.
  5. [UPCOMING] QV File Reports: checks available files on server. As soon as ICOS allows.

Site Reports


SRP (LH, fluxes): Station was offline, no new fluxes for the moment.
Tech-SRP (IB): –


SRP (KF): The farmer cut the grass at both parcels (27th April 2018) and we went biomass sampling for the first time this year (3rd May 2018). Iris got to know the station and the responsible people on site. The IRGA is running nicely, however the QCL reparation by Aerodyne will take more time.
Tech-SRP (PL): –


SRP (MG, LH): Due to factory calibration of the LI-7200, we have a large gap in our eddy flux calculations. Data are missing between 7 Mar 2018 9:00 and 30 Apr 2018 15:00 (> 54 days).
Tech-SRP (PM): –


SRP (LH): Site is running very well, meteo data look good, same for fluxes.
Tech-SRP (PL): –


SRP (EPL): –
Tech-SRP (TB): –


SRP (RM): Everything OK until 15.5.2018, since then the station is down. Techs are informed and working on it.
Tech-SRP (TB): –


You are here: Short Report - QA/QC Meeting May 2018


Last Updated on 14 Feb 2024 12:14

Leave a Reply

You must be logged in to post a comment.